Vcloud - Vground = 1.1 x 108 V, with the cloud being at the higher potential. What is the change in an electron's potential energy when the electron moves from the ground to the cloud?
1
Expert's answer
2020-09-15T10:05:28-0400

The change in an electron's potential energy is equivalent to the work made by electron in electric field.

"\Delta U = W = e \Delta V = 1.6 \cdot 10^{-19} \cdot 1.1 \cdot 10^8 = 1.76 \cdot 10^{-11}" J.